About My Writing Intensive Division of Classical Algebra (Mat 326Z)

Call No. 2697

William F. Hammond

Fall Semester, 2002

The University’s upper division writing requirement resides under the concept of disciplineoriented writing. In this course the writing exercises will involve (i) presenting problem solutions, (ii) writing mathematical arguments, and (iii) writing simple computer programs. It will also be necessary to present some of these assignments on the web although it will not be required that the web address be advertised. Students in the writing intensive section must complete all assignments in a satisfactory manner. (It will not be possible to elude any of these requirements.)

Students enrolled in the writing intensive version of the course, Math 326Z, will not receive a grade higher than C- without satisfactory use of mathematical English on tests and quizzes and particularly on the submitted assignments.

Satisfactory use of mathematical English for this course requires satisfactory use of general English prose at a level that possibly alone suffices for handling the upper division writing requirement for courses in some of the other disciplines.

Students who have yet to satisfy the lower division writing intensive requirement may not register for my division of MAT 326Z.

Past experience has shown that some mathematics students who have attempted to satisfy the University’s upper division writing intensive requirement in this course have had serious trouble in the course because of difficulties with their writing.

For this reason the following procedure will be used to control pre-registration admission to my writing intensive division:

1. Register for my regular division (call no. 2695) of Classical Algebra (Mat 326). No SKN should be required unless the section fills. (If the section fills and you wish to try to gain entrance in my division of Math 326, writing intensive or not, contact me for entry on a waiting list.)

2. Prepare a formal essay, 250 to 500 words in length, describing your interests in and experiences with mathematics. Attach an appendix to the essay giving

(a) the name under which you are registered at the University.

(b) your student identification number.

(c) the total number of graduation credits you expect to have completed by the end of the current semester.

(d) the course that was used to satisfy your lower division writing intensive requirement.

(e) an email address or, if none is available, a regular address where you can be reached after the end of the semester.

3. Present the essay at the office of the Department of Mathematics and Statistics, Earth Science 110, marked for delivery to me.